OpenStudy (anonymous):

Complete the ordered pair so that it is a solution for the equation 3x-4y=24 (2,?)

OpenStudy (anonymous):

thats wrong..

hero (hero):

No it isn't

OpenStudy (anonymous):

its suppose to equal 24

hero (hero):

3x-4y=24 3(2)-4y = 24 6 - 4y = 24 -4y = 18 y = 18/-4 y = -4.5

hero (hero):

I stand corrected

hero (hero):

(2,-4.5)

OpenStudy (anonymous):

I check it over and its right, thankyou! would would be the answer for this one? (?,6)

hero (hero):

I could show you how to solve these in real time. Would you like to see it?

OpenStudy (anonymous):

yea, sure

hero (hero):

Okay, but I have certain platforms that I use. I hope you'll be okay with them. I use Skype, MS One Note, and Zoho Assist

OpenStudy (anonymous):

yea, thats fine

hero (hero):

What is your skype?

OpenStudy (anonymous):

don't have one, just show me how to do it on here please